OpenStudy (anonymous):

The probability that Roger wins a tennis tournament is 0.75 (event A), and the probability that Stephan wins a golf tournament is 0.64 (event B). The probability that both Roger and Stephan win their respective tournaments is 0.48.

OpenStudy (anonymous):

Events A and B are independent because P(A and B) = P(A) P(B). Events A and B are independent because P(A and B) = P(A) + P(B). Events A and B are dependent because P(A and B) = P(A) P(B). Events A and B are dependent because P(A and B) = P(A) + P(B). Events A and B are dependent because P(A and B) P(A) P(B).

OpenStudy (kirbykirby):

Two events are independent if: \(P(A \text{ and } B)=P(A)P(B)\) You are given that \(P(A \text { and } B)=0.48\), so you can find if their independent if P(A)*P(B) is indeed equal to 0.48